### v2

# wd1: Weighted Policy Optimization for Reasoning in Diffusion Language Models

Abstract
Improving the reasoning capabilities of diffusion-based large language models (dLLMs) through reinforcement learning (RL) remains an open problem. The intractability of dLLMs likelihood function necessitates approximating the current, old, and reference policy likelihoods at each policy optimization step. This reliance introduces additional computational overhead, and can lead to large variance and estimation error in RL objective – particularly in computing the policy ratio for importance sampling. To mitigate these issues, we introduce wd1, a novel ratio-free policy optimization approach that reformulates the RL objective as a weighted log-likelihood, requiring only a single approximation for the current parametrized policy likelihood. We formally show that our proposed method can be interpreted as energy-guided discrete diffusion training combined with negative sample unlearning, thereby confirming its theoretical soundness. In experiments on LLaDA-8B model, wd1 outperforms diffusion-based GRPO (d1) while requiring lower computational cost, achieving up to a $59\%$ improvement in accuracy. Furthermore, we extend wd1 to denoising-stepwise weighted policy optimization (wd1++), achieving state-of-the-art math performance of $44.2\%$ on MATH500 and $84.5\%$ on GSM8K with only 20 RL training steps.
1 Introduction
Diffusion-based large language models (dLLMs) have recently gained attention as promising alternatives to autoregressive (AR) models for language modelling tasks (Nie et al., 2025b; Ou et al., 2025; Yang et al., 2025). Unlike AR models, which generate tokens sequentially, dLLMs iteratively refine entire response sequences through a denoising process. A primary advantage of this approach is the significantly improved inference efficiency. Notably, recent closed models such as Mercury (Labs et al., 2025) and Gemini Diffusion achieve over an order of magnitude speed-up in generation compared to AR models, while maintaining comparable generation quality. Furthermore, open-weight dLLMs demonstrate competitive performance on standard language benchmarks, with smaller models (Lou et al., 2024; Ou et al., 2025; Nie et al., 2024) achieving parity with equivalently sized AR baselines, and larger-scale models such as LLaDA-8B (Zhu et al., 2025a) and Dream-7B (Ye et al., 2025) extending this trend at scale. While dLLMs demonstrate strong performance in text generation, it remains an open and important question how best to fine-tune dLLMs using RL – a paradigm that has proven highly effective in alignment and improving reasoning capabilities of AR models (Ouyang et al., 2022; Shao et al., 2024).
A key challenge in applying reinforcement learning (RL) to dLLMs is the intractability of their likelihood functions (Zhao et al., 2025; Yang et al., 2025), which necessitates approximation for policy optimization. Applying approximated log-likelihood for diffusion-based GRPO (Shao et al., 2024; Zhao et al., 2025) can exponentially amplify the approximation error and lead to large variance when computing the policy ratio for importance sampling. Moreover, GRPO requires the estimated likelihoods of the current, old, and reference policies at every training step, leading to significant computational overhead. These issues can be further exacerbated as the completion length and the number of diffusion steps increase.
To address these challenges, we propose wd1, a policy optimization approach with weighted log-likelihood objective for dLLMs. Crucially, this objective dispenses with explicit policy ratios and relies on a single likelihood approximation, thereby avoiding the potentially large bias and variance in policy ratio, and reducing the computational overhead. Our principal contributions are:
  • We propose a novel reinforcement learning method for dLLMs, termed wd1, which formulates the objective as a weighted log-likelihood of outcome sequence, derived from reverse-KL regularized policy optimization. The weight, defined as $(-w^{+}+w^{-})$ and dependent on the advantage $A$, balances two terms: $w^{+}\propto\exp(A)$ increases the probability of higher-advantage completions, while $w^{-}\propto\exp(-A)$ decreases the probability of lower-advantage ones. Together, this mechanism amplifies beneficial outcomes meanwhile actively reducing detrimental ones.
  • We prove that our proposed RL method for dLLMs can be interpreted as jointly training an energy-guided discrete diffusion model—guided by the advantage function—and unlearning low-advantage data, thereby steering generation toward higher-advantage completions.
  • We conduct experiment with LLaDA-8B-Instruct model (Nie et al., 2025a). Compared to the baseline method d1 (Zhao et al., 2025), our method wd1 achieves 76.4% on Sudoku (Arel, 2025) (+58.8% over d1) and 51.2% on Countdown (Pan et al., 2025) (+16% over d1), without requiring supervised fine-tuning (SFT), and with significantly less computational burden in RL training.
  • We further extend our method to leverage intermediate completions generated in the decoding process, which we call wd1++. The extended method surpasses several concurrent RL for dLLMs methods, achieving state-of-the-art performance 44.2% on MATH500 and 84.5% on GSM8K with only 20 training steps, and $10\times$ fewer rollouts compared to the baseline methods.
2 Preliminaries
We denote the generation policy of diffusion-based Large Language Models (dLLMs) by $\pi_{\theta}$. Denote prompt $q\in\mathcal{D}$, and completions $o\in\mathcal{O}$. Notably, the reward function denoted by $R(q,o)$ in this work is not limited to verifiers. We use superscript $k$ to indicate the $k$-th token of completion: $o^{k}$ or $x_{0}^{k}$.
2.1 Diffusion Large Language Models
The most promising discrete diffusion for language modeling is masked diffusion, which gradually corrupts the text sequence with a mask token (Sahoo et al., 2024; Ou et al., 2025; Shi et al., 2025; Lou et al., 2024). Let $t\in[0,1]$ denote the diffusion timestep, and $x_{t}$ as the masked sequence at step $t$. The fully denoised sequence (i.e., the completion $o$) is represented by $x_{0}$, and the forward process ($p_{t|0}(x_{t}\mid x_{0})$) is formulated as a continuous-time Markov chain. The transition kernel $\mathbf{Q}_{t}$ is absorbing (Austin et al., 2023; Campbell et al., 2022), meaning that at time $t$, $\mathbf{Q}_{t}=\sigma(t)\mathbf{Q}^{\text{absorb}}$, where $\sigma$ is a decreasing scalar noise schedule and $\mathbf{Q}^{\text{absorb}}$ is a constant matrix (See Definition 2).
This work aims to apply reinforcement learning to fine-tune masked discrete diffusion models such as LLaDA (Ou et al., 2025; Zhu et al., 2025a), which models the clean data distribution conditional on masked sequence as $\pi_{\theta}(x_{0}^{k}\mid x_{t})$. The negative Evidence Lower Bound (ELBO) is reduced to a simple objective called Denoising Cross Entropy (DCE) (Ou et al., 2025): $\forall x_{0}\sim p_{\text{data}}$,
$\mathcal{L}_{\text{DCE}}(x_{0})=-\mathbb{E}_{t\sim\mathcal{U}[0,1],\ x_{t}\sim p_{t|0}(x_{t}|x_{0})}\left[\frac{1}{t}\sum_{k=1}^{K}\mathbf{1}(x_{t}^{k}=[\texttt{mask}])\log\pi_{\theta}(x_{0}^{k}\mid x_{t})\right],$ (1)
where $K$ is the length of the sequence and $x_{0}^{k}$ is the $k$-th token of $x_{0}$. Specifically, intermediate steps $t$ are sampled from uniform distribution, and masked sequence is sampled following the predefined forward process $p_{t|0}(x_{t}\mid x_{0})$. DCE can be used to approximate the marginal likelihood $\log\pi_{\theta}(x_{0})$ for supervised fine-tuning and reinforcement learning (Nie et al., 2025a; Yang et al., 2025).
2.2 Existing Policy Optimization Methods
The base method of current prevailing RL fine-tuning algorithms is Trust Region Policy Optimization (TRPO) (Schulman et al., 2015), in which forward KL divergence is applied to restrict the update:
[…]
where $A^{\pi_{\text{old}}}$ is the advantage function, $q$ and $o$ are denoted as the prompt and (clean) response, respectively. Proposition˜1 (Appendix˜A) demonstrates the monotonic policy improvement of TRPO.
PPO then extends the soft constraint (KL penalty) to clipping policy ratio $\pi_{\theta}(\cdot|q)/\pi_{\text{old}}(\cdot|q)$ and employing pessimism for policy update, further employed in fine-tuning (Ouyang et al., 2022) with additional reverse-KL regularization w.r.t. the reference policy $\pi_{\text{ref}}$. Group Relative Policy Optimization (GRPO) (Shao et al., 2024) simplifies PPO by sampling a group of completions $\{o_{i}\}_{i=1}^{G}$ and approximating their advantage with their normalized rewards. This advantage is corrected by subtracting the mean reward across the group (Liu et al., 2025): $\hat{A}_{i}=R(q,o_{i})-\texttt{mean}\big(R(q,o_{1:G})\big),$ which we refer to as the group-relative advantage.
2.3 Policy Optimization for dLLMs
Adapting GRPO to diffusion-based large language models (dLLMs) presents notable challenges, since dLLMs generate outputs via a non-autoregressive, iterative denoising process, making the computation of $\log\pi_{\theta}(o|q)$ intractable and necessitating approximation for policy optimization.
Existing works by Nie et al. (2025a); Yang et al. (2025) employ ELBO for per-token log-likelihood approximation following DCE: $\phi^{\pi}(x_{0}^{k})=\mathbb{E}_{t\in\mathcal{U}[0,1]}[w\cdot\mathbf{1}[x_{t}^{k}=\mathtt{mask}]\log\pi(x_{0}^{k}|x_{t},q)]$, where $w=1/t$ in DCE and $w=1$ in UniGRPO (Yang et al., 2025). However, an accurate estimation requires a large sample size of $t$, resulting in inefficiency for online RL. A biased but efficient method is introduced in d1 (Zhao et al., 2025), requiring only sample at $t=1$: $\phi^{\pi}(x_{0}^{k})=\log\pi(x_{0}^{k}|x_{1},q^{\prime})$, where prompt $q^{\prime}$ is randomly masked, $x_{1}$ is fully masked response.
In diffusion-based GRPO (Zhao et al., 2025; Yang et al., 2025), policy ratio is then computed using the approximated log-likelihoods: $r^{k}_{i}(\theta)=\pi_{\theta}(o^{k}_{i})/\pi_{\text{old}}(o^{k}_{i})\approx\exp\big(\phi^{\pi_{\theta}}(o_{i}^{k})-\phi^{\pi_{\text{old}}}(o_{i}^{k})\big)$ for importance sampling in estimating the objective of GRPO:
$\mathbb{E}_{\begin{subarray}{c}q\sim\mathcal{D},\\ \,o_{1:G}\sim\pi_{\text{old}}(\cdot\mid q)\end{subarray}}\bigg[\frac{1}{GK}\sum_{i=1}^{G}\sum_{k=1}^{K}\min\big(r^{k}_{i}(\theta)\hat{A}_{i},\text{clip}(r^{k}_{i}(\theta),1\pm\epsilon)\hat{A}_{i}\big)-\beta D_{\text{KL}}\big(\pi_{\theta}(\cdot)\,\|\,\pi_{\text{ref}}(\cdot)\big)\bigg].$ (3)
Figure 1: Example policy ratio value $r_{i}^{k}$ computed using ELBO and approximated likelihood in d1 on GSM8K after a policy update. Ratio’s unclipped interval is $1-\epsilon,1+\epsilon]$, where $\epsilon=0.5$. ELBO-based likelihood approximation yields high-variance ratio estimates; d1 induces a biased ratio that can deviate substantially from ELBO. Both methods suffer from efficiently and accurately compute ratios.
However, existing approaches are hampered by their reliance on extensive likelihood approximation to compute the policy ratio. In current diffusion-based GRPO methods, the ratio is computed as $r_{i}^{k}\approx\exp\big(\phi^{\pi_{\theta}}(o_{i}^{k})-\phi^{\pi_{\text{old}}}(o_{i}^{k})\big)$ so approximation errors in likelihood can be exponentially amplified. As formally shown in Appendix A.1, the resulting error in the estimated RL objective becomes more severe when less accurate log-likelihood approximations are used, such as in d1, or ELBO used in DCE and UniGRPO when the Monte Carlo sample size $t$ is small.
[…]
3 wd1: Weighted Policy Optimization for dLLMs
In this section, we introduce wd1, a novel RL algorithm that eliminates the need for approximating the likelihood (policy) ratios for importance sampling, aiming to reduce the computational burden, and the variance and approximation error in computing the RL objective. We further extend our method to wd1++ by applying denoising-stepwise policy optimization.
3.1 Reinforcement Learning as Weighted Log-Likelihood Maximization
Prevailing RL methods are based on constrained policy optimization (Belousov and Peters, 2017), penalizing the deviation of current policy $\pi_{\theta}(\cdot|q)$ from the old policy $\pi_{\text{old}}(\cdot|q)$. TRPO objective (Equation˜2) applies a forward-KL penalty. We instead adopt reverse-KL penalty augmented with the reference policy regularization $D_{\text{KL}}(\pi_{\theta}(\cdot|q)\,\|\,\pi_{{\text{ref}}}(\cdot|q))$:
[…]
Note that the monotonic improvement guarantee still holds when using reverse-KL penalty, as we show in Theorem 2. From the method of Lagrange multipliers, the solution to Equation˜4 has the following form (Peng et al., 2019; Rafailov et al., 2023):
$\pi^{*}(\cdot|q)\propto\pi_{\text{old}}(\cdot|q)^{\lambda/(\lambda+\beta)}\cdot\pi_{{\text{ref}}}(\cdot|q)^{\beta/(\lambda+\beta)}\cdot\exp\bigg(\frac{A^{\pi_{\text{old}}}(q,\cdot)}{\lambda+\beta}\bigg).$ (5)
As the analytic form of the optimal policy $\pi^{*}$ is known, we can train our policy by directly minimizing $D_{\text{KL}}(\pi^{*}(\cdot|q)\,\|\,\pi_{\theta}(\cdot|q))$. This minimization can be expressed as the following weighted log-likelihood (WLL) loss, where the weights $\propto\exp\big(\psi A^{\pi_{\text{old}}}\big)$, $\psi=\tfrac{1}{\lambda+\eta}$ and the samples are obtained from the geometric mixture policy $\pi_{\text{old}}^{\text{ref}}(\cdot|q)\propto\pi_{\text{old}}(\cdot|q)^{\lambda/(\lambda+\beta)}\cdot\pi_{{\text{ref}}}(\cdot|q)^{\beta/(\lambda+\beta)}$ (See Proposition˜2): $\forall q\sim\mathcal{D}$,
$\mathcal{L}_{\text{WLL}}(\theta)$ $=\mathbb{E}_{o\sim\pi^{\text{ref}}_{\text{old}}(\cdot|q)}\Big[-\exp\big(\psi A^{\pi_{\text{old}}}(q,o)\big)\cdot\log\pi_{\theta}(o|q)\Big]$ (6)
$\approx\mathbb{E}_{\{o_{i}\}_{i=1}^{G}\sim\pi^{\text{ref}}_{\text{old}}(\cdot|q)}\left[\frac{1}{G}\sum_{i=1}^{G}-\frac{\exp\big(\psi\hat{A}_{i}\big)}{\sum_{j=1}^{G}\exp\big(\psi\hat{A}_{j}\big)}\log\pi_{\theta}(o_{i}|q)\right].$ (7)
As shown in Equation˜7, we approximate the advantage function using the group-relative advantage $\hat{A}$ and normalize the weights, thereby limiting their magnitude and reducing variance in loss computation. Notably, dividing by the normalization constant does not affect the solution, since it is independent of the completions. The resulting objective does not involve ratio $\pi_{\theta}(\cdot|q)/\pi_{\text{old}}(\cdot|q)$ for importance sampling or $\pi_{\theta}(\cdot|q)/\pi_{\text{ref}}(\cdot|q)$ for regularization, successfully avoiding the potential amplification of log-likelihood approximation error and large variance in diffusion GRPO.
Although the objective $\mathcal{L}_{\text{WLL}}(\theta)$ in Equation˜7 avoids the likelihood ratio estimation, it has two limitations. First, the algorithm is not fully utilizing all the completions. Due to the exponential form of the weighting, completions with relatively low advantage – referred to as negative samples – may receive vanishingly small weights, and do not contribute to learning. Second, due to the likelihood-maximization property of WLL, the likelihoods of all sampled completions are increased, even for negative samples. This issue is exacerbated in scenarios where all completions attain identical but low rewards (e.g. 0), thus all weights become equal and the likelihoods of these suboptimal samples are nonetheless reinforced.
3.2 wd1: Fully Utilizing Completions
We propose wd1, an improved weighted log-likelihood objective that explicitly reinforces positive samples and penalizes negative samples:
$\mathcal{L}_{\textit{wd1}}(\theta)=\mathbb{E}_{q\sim\mathcal{D},\{o_{i}\}_{i=1}^{G}\sim\pi_{\text{old}}^{\text{ref}}(\cdot|q)}\Big[\frac{1}{G}\sum_{i=1}^{G}\big(-w^{+}(q,o_{i})+w^{-}(q,o_{i})\big)\cdot\log\pi_{\theta}(o_{i}|q)\Big],$ (8)
where the weights are based on group-relative (GRPO) advantage and are further normalized to avoid overly imbalanced weight $\hat{A}_{i}=R(q,o_{i})-\texttt{mean}(R(q,o_{1:G}))$:
$w^{+}(q,o_{i})=\frac{\exp\big(\psi\hat{A}_{i}\big)}{\sum_{j=1}^{G}\exp\big(\psi\hat{A}_{j}\big)},\quad\ w^{-}(q,o_{i})=\frac{\exp\big(-\psi\hat{A}_{i}\big)}{\sum_{j=1}^{G}\exp\big(-\psi\hat{A}_{j}\big)}.$ (9)
wd1 objective balances positive and negative samples through a complementary penalty term, $w^{-}(q,o_{i})\,\log\pi_{\theta}(o_{i}|q)$, which minimizes the likelihood of low-advantage completions. This penalty induces negative gradients, thereby accelerating divergence from undesirable completions. Moreover, in the extreme case where all completions exhibit identical advantages, the optimization naturally halts since $w^{+}=w^{-}$, thereby addressing the concern on increasing likelihood of negative samples proposed in Sec 3.1. We demonstrate the effectiveness of this combination via ablations in C.2.
Our method wd1, a simple ratio-free policy optimization based on weighted log-likelihood objective for dLLMs, is formally presented in Algorithm˜1. We first obtain $G$ completions $\{o\}_{i=1}^{G}$ sampled from geometric mixture $\pi_{\text{old}}^{\text{ref}}(\cdot|q)\propto\pi_{\text{old}}(\cdot|q)^{\lambda/(\lambda+\beta)}\cdot\pi_{{\text{ref}}}(\cdot|q)^{\beta/(\lambda+\beta)}$ (line 5). Since the base model LLaDA parametrizes the clean token prediction $\pi^{\text{ref}}_{\text{old}}(x_{0}^{k}|x_{t},q)$ for denoising, we approximate $\log\pi^{\text{ref}}_{\text{old}}(x_{0}^{k}|x_{t},q)\approx\lambda\log\pi_{\text{old}}(x_{0}^{k}|x_{t},q)+\beta\log\pi_{\text{ref}}(x_{0}^{k}|x_{t},q)$ as the logits of the denoising distribution at each step $t$. We then use the samples to compute weights in Equation˜9 (line 6). In weights computing, we leverage completions from all groups to estimate the normalization constant, in order to restrict the the gradient norm and stabilize the training. Finally in line 8, we approximate the log-likelihood of completions, and compute objectives for policy update. Likelihood approximation in d1 (Zhao et al., 2025) is directly applicable to wd1: $\log\pi_{\theta}(x_{0}|q)\approx\sum_{k}\log\pi_{\theta}(x_{0}^{k}|x_{1},q^{\prime})$, where $q^{\prime}$ is randomly masked from prompt $q$ at every gradient step.
3.3 wd1++: Stepwise Weighted Policy Optimization
The decoding process in dLLMs relies on confidence-based remasking (Wang et al., 2025b). At each denoising step $l\in\{1,\cdots,L\}$ in decoding, clean data is predicted conditional on the masked sequence $x_{l}$ and then tokens with low-confidence are re-masked for further denoising, which construct a refinement process. Since current diffusion RL methods only use the final predicted clean completion for training, there are bunch of clean completions in the intermediate denoising steps remaining unused.
To leverage intermediate clean completions, we extend our weighted log-likelihood objective to a step-wise formulation based on DCE, which we term wd1++. In wd1 (as well as in GRPO), a group of completions $\{o_{i}\}_{i=1}^{G}$ is sampled for policy optimization. In wd1++, we expand this group to $\{O_{i}\}_{i=1}^{G}$, where $O_{i}=\{x_{0|l}\mid x_{0|l}\sim\pi_{\text{old}}^{\text{ref}}(\cdot\mid x_{t},q),\;x_{0|L}=o_{i},\;l=1,\dots,L\},$ which contains all generated completions during the decoding process, including intermediate ones. The expanded group of completions is then used to estimate both the advantage function and the corresponding weights. The resulting loss objective is defined as:
$\mathcal{L}_{\textit{wd1}++}(\theta)=\mathbb{E}_{\begin{subarray}{c}q\sim\mathcal{D},\\ \{O_{i}\}_{i=1}^{G}\sim\pi_{\text{old}}^{\text{ref}}(\cdot|q)\\ l\in\text{Unif}\{1,\cdots,L\}\end{subarray}}\big[\tfrac{L}{Gl}\sum_{i=1}^{G}\sum_{x_{0|l}\in O_{i}}\big(-w^{+}(q,x_{0|l})+w^{-}(q,x_{0|l})\big)\cdot\log\pi_{\theta}(x_{0|l}|x_{l},q)\big].$ (10)
Algorithm 1 wd1: Weighted Policy Optimization for dLLMs
Require: Reference model $\pi_{\text{ref}}$, prompt distribution $\mathcal{D}$, group size $G$, reward function $R$, dLLM $\pi_{\theta}$, regularization hyperparameters $\lambda$ and $\beta$
1: Initialize $\pi_{\theta}\leftarrow\pi_{\text{ref}}$
2:while not converged do
3:  $\pi_{{\text{old}}}\leftarrow\pi_{\theta}$
4:  Sample prompt $q\sim\mathcal{D}$ and $G$ completions $o_{i}\sim\pi_{{\text{old}}}(\cdot\mid q),\forall i\in[G]$
5:  Compute advantage $\hat{A}_{i}=R(q,o_{i})-\texttt{mean}(R(q,o_{1:G}))$, $\forall i\in[G]$
6:  Compute weights $w^{+}$ and $w^{-}$ in Equation˜9, $\forall i\in[G]$
7:  for gradient update iterations $n=1,2,\dots,\mu$ do
8:   Compute approximated log-likelihood $\log\pi_{\theta}(o_{i}|q)$
9:   Compute objective $\mathcal{L}_{\textit{wd1}}(\theta)$ in Equation˜8 or Equation˜10 and update $\theta$
10:  end for
11:end while
12:return $\pi_{\theta}$
4 Theoretical Insights: Energy-Guided Diffusion Sampling
In this section, we present a novel theoretical interpretation of policy optimization for dLLMs. We prove that the advantage-weighted log-likelihood objective (wd1) for dLLMs can be viewed as energy-guided discrete diffusion training combined with negative sample unlearning.
Sampling from the solution policy of the reverse-KL policy optimization, as described in Equation˜5, can be interpreted as energy-guided sampling, where the energy function $\mathcal{E}(q,\cdot)=-{A^{\pi_{\text{old}}}(q,\cdot)}$. Equation˜5 defines the marginal distribution of the clean data ($x_{0}=o$) which we denote as $p^{*}_{0}(x_{0})$ [^1] To adapt to the setting of diffusion, we use $x_{t}$ to denote the (masked) completions, and omit the prompt $q$.. To obtain the guidance at intermediate time steps $t>0$, we define the forward diffusion process for the target diffusion policy $\pi^{*}$ as following.
Definition 1.
The forward diffusion process of the target policy ($\pi^{*}$) satisfies $p^{*}_{t|0}(x_{t}|x_{0})=p_{t|0}(x_{t}|x_{0})$, where $p_{t|0}$ is the forward process of old diffusion policy $\pi_{\text{old}}$.
Since the reference diffusion policy is the initial policy, three policies have identical forward diffusion process, being $p^{*}_{t|0}(x_{t}|x_{0})=p_{t|0}(x_{t}|x_{0})=p^{\text{ref}}_{t|0}(x_{t}|x_{0})$, and thus, $p^{*}_{t|0}(x_{t}|x_{0})=p^{\prime}_{t|0}(x_{t}|x_{0})$, where $p^{\prime}$ is the geometric mixture diffusion $p^{\prime}_{t|0}(x_{t}|x_{0})\propto p_{t|0}(x_{t}|x_{0})^{\lambda}p^{\text{ref}}_{t|0}(x_{t}|x_{0})^{\beta}$. We can then obtain the energy guidance at all time step $t$.
Lemma 1 (Intermediate Energy Guidance on Discrete Diffusion).
The marginal probability distribution of the masked responses ($x_{t}$) in the diffusion process satisfies $p^{*}_{t}(x_{t})=p^{\prime}_{t}(x_{t})\cdot\exp\big(A_{t}(x_{t})\big)/Z_{t}$, which induces an energy-guided discrete diffusion:
$p^{*}_{0|t}(x_{0}|x_{t})\propto p^{\prime}_{0|t}(x_{0}|x_{t})\cdot\exp(A(x_{0})-A_{t}(x_{t})),$ (11)
where $A_{t}(x_{t})=-\log\mathbb{E}_{x_{0}\sim p^{\prime}_{0|t}(\cdot|x_{t})}[\exp\big(A(x_{0})\big)]$ is intermediate energy function for $t>0$, and $A(\cdot)$ is advantage function (Proof in ˜Lemma).
The guidance provided in Lemma˜1 demonstrates that it directs the sampling process toward generating completions that exhibit higher advantage values. However, conducting training-free guided sampling following Equation˜11 requires estimating the posterior mean of the exponential of advantage (Lu et al., 2023). Rather than relying on such estimation, we instead aim to find the training objective to directly approximate the target guided diffusion model.
Since existing masked dLLMs parametrize the concrete score (Meng et al., 2022), to apply the energy guidance, we aim to directly approximate target guided concrete score. Denote $x_{t}=(x_{t}^{1},\cdots,x_{t}^{d})$ and $\hat{x}_{t}$ is identical to $x_{t}$ except the $i$-th token is unmasked (i.e. $x_{t}^{i}=[M]$ and $\hat{x}_{t}^{i}\neq[M]$). Concrete score is defined as the marginal probability ratio between $\hat{x}_{t}$ and $x_{t}$:
$s(x_{t},t)\overset{\text{def}}{=}\frac{p(x_{t}^{1},\cdots,\hat{x}_{t}^{i},\cdots,x_{t}^{d})}{p(x_{t}^{1},\cdots,x_{t}^{i},\cdots,x_{t}^{d})}.$ (12)
We prove that the training objective to approximate the guided concrete score can be simplified as a weighted Denoising Concrete Score Matching (D-CSM) (Meng et al., 2022):
Theorem 1.
The model $s_{\theta}$ approximates the concrete score of the energy-guided discrete diffusion $p^{*}$ when the following loss objective is minimized. This objective is in a form of advantage-weighted Denoising Concrete Score Matching, which we call AW-D-CSM:
$\mathcal{L}_{\text{AW-D-CSM}}=$ $\mathbb{E}_{x_{0}\sim p^{\prime}_{0}(\cdot)}\bigg[\underbrace{\exp\big(A(x_{0})\big)}_{\text{Advantage Weight}}\cdot\underbrace{\mathbb{E}_{t\sim[0,T],p^{\prime}_{t|0}(x_{t}|x_{0})}[\|{s_{\theta}(x_{t},t)}-\frac{p^{\prime}_{0}(\hat{x}_{t}|x_{0})}{p^{\prime}_{0}(x_{t}|x_{0})}\|_{2}^{2}]}_{\mathcal{L}_{\text{D-CSM}}(x_{0})}\bigg].$ (13)
We provide the proof in Appendix Theorem. Additionally, D-CSM is an approximation of CSM (Meng et al., 2022), which is equivalent to Denoising score entropy (DSE) (Lou et al., 2024). For all $x_{0}$, it is satisfied up to multiplying a constant that $\mathcal{L}_{\text{D-CSM}}(x_{0})\Leftrightarrow\mathcal{L}_{\text{CSM}}(x_{0})\Leftrightarrow\mathcal{L}_{\text{DSE}}(x_{0})\Leftrightarrow\mathcal{L}_{\text{DCE}}(x_{0})$ (Ou et al., 2025). Therefore, AW-D-CSM can then be applied for both SEDD (Lou et al., 2024) and RADD (Ou et al., 2025) model such as LLaDA. Denote $p_{\theta}$ as the concrete score reparametrized model, AW-D-CSM can be converted to a weighted denoising cross-entropy loss (AW-DCE):
$\mathcal{L}_{\text{AW-DCE}}=$ $\mathbb{E}_{x_{0}\sim p^{\prime}_{0}(\cdot)}\bigg[\exp\big(A(x_{0})\big)\cdot\mathbb{E}_{t\sim[0,T],p^{\prime}_{t|0}(x_{t}|x_{0})}\big[\sum_{x^{i}_{t}=[\texttt{mask}]}-\frac{1}{t}\log p_{\theta}(x_{0}^{i}|x^{\text{UM}}_{t})\big]\bigg].$ (14)
DSE and DCE objectives both can be used for likelihood approximation in fine-tuning (Ou et al., 2025; Nie et al., 2025a; Yang et al., 2025) since they can serve as negative ELBO (Lou et al., 2024; Shi et al., 2025). Thus, the advantage-weighted objective AW-DCE (or AW-DSE) used to learn energy-guided score is in a weighted log-likelihood form as in wd1 with only $w^{+}$ (i.e. WLL loss in Equation˜6), which contributes to our main theoretical findings:
Remark 1.
In the context of applying RL to masked discrete diffusion, the advantage-weighted log-likelihood (WLL) objective (Equation˜6) induced by reverse-KL policy optimization, is equivalent to the objective of training energy-guided diffusion models, where the energy function is the negative advantage. Formally, $\mathcal{L}_{\text{WLL}}\Leftrightarrow\mathcal{L}_{\text{AW-DCE}}$ when DCE is used for likelihood approximation.
Remark 2.
Additionally, based on DCE likelihood, the additional penalty term on negative samples used to extend WLL to wd1 loss can be viewed as applying data unlearning by minimizing the ELBO (Alberti et al., 2025), where the data $\{x^{-}_{0}\}$ (negative samples) has probability distribution $p_{\text{data}}(x^{-}_{0})\propto p^{\prime}_{0}(x^{-}_{0})\exp(-A(q,x^{-}_{0}))$, which corresponds to a Boltzmann distribution that places higher probability mass on regions with lower advantage values (more details in Section˜D.1).
5 Experiments
In this section, we empirically validate the following key advantages of our approach:
  1. i)
    Improved reasoning capabilities than existing methods on popular reasoning benchmarks;
  2. ii)
    reduced computational burden, as reflected by decreased runtime, lower FLOPs and numbers of function evaluations (NFEs) per training step, number of training steps and rollouts; and
  3. iii)
    marked performance gains attributable to the incorporation of samples with low-advantage.
To evaluate our approach, we next detail the experimental setup and implementation.
Experimental Setup. We perform reinforcement learning (RL) fine-tuning on the LLaDA-8B-Instruct model (Nie et al., 2025a) with Low-Rank Adaptation (LoRA) on: GSM8k (Cobbe et al., 2021), MATH (Lightman et al., 2023), Sudoku (Arel, 2025), and Countdown (Pan et al., 2025). As for decoding, we follow the default strategy Mounier and Idehpour (2025); Arriola et al. (2025); Wang et al. (2025b). Our main baseline is d1 (Zhao et al., 2025), the first RL method developed for masked diffusion LLMs (dLLMs). We reproduce the baseline methods Diffu-GRPO, which applies diffusion-based GRPO training directly to the LLaDA base model, and d1, which performs SFT before applying Diffu-GRPO. We use s1K  (Muennighoff et al., 2025) data for SFT in d1. We also compare with SDPO (Han et al., 2025), TCR (Wang et al., 2025d), and MDPO (He et al., 2025) on benchmarks GSM8K and MATH500. MDPO is reproduced based on the official implementation and the training dataset (He et al., 2025).
Implementation. As for wd1, we conduct training on the same dataset as in d1 (Zhao et al., 2025): training splits on GSM8k and MATH, and the dataset splits provided by Zhao et al. (2025) on Sudoku and Countdown. In our implementation of wd1, we apply the same likelihood approximation method as d1. The hyperparameters used in our method and our reproduction of d1 are listed in Table˜6 and Table˜5. As for wd1++, we train on dataset provided by (He et al., 2025), which is sampled from OpenR1 dataset (Face, 2025). Since previous works (Yu et al., 2025) have demonstrated that the reference policy is empirically unnecessary, we set $\beta=0$ and $\lambda=1$ to eliminate $\pi_{\text{ref}}$ in practice. We report results using zero-shot and pass@1 evaluation on sequence lengths of 256 and 512 tokens.
Table 1: Test Accuracy (%) of wd1 and d1. We reproduce d1 and vary completion length. Our approach without SFT, demonstrates particularly higher accuracy on Sudoku [^3] In the technical report version of this work, our method achieved scores of 25.2 and 24.2 on Sudoku after 5K training steps. In this paper, we extend the training to 12.5K steps, and wd1 results in improved performance.and Countdown.
Model / Gen Len Sudoku Countdown GSM8K MATH500
256 512 256 512 256 512 256 512
LLaDA-8B-Instruct 6.7 5.5 19.5 16.0 76.7 78.2 32.4 36.2
+ diffu-GRPO 16.1 11.7 27.0 34.0 80.7 79.1 34.4 39.0
+ SFT + diffu-GRPO (d1) 17.6 16.2 25.8 35.2 78.2 82.0 34.4 38.0
+ wd1 76.4 62.8 51.2 46.1 80.8 82.3 34.4 39.0
[…]
Method SFT RL Training
Time Cost Time Cost FLOPs NFEs for Likelihood
d1 2.01 hrs 103.5 sec/step $9.922\times 10^{15}$/step $(\mu+2)$/step
wd1 0 hrs 81.16 sec/step $8.887\times 10^{15}$/step $\mu$/step
5.1 Main Results
Superior Reasoning Ability. In Table˜1, we observe that wd1, even without supervised fine-tuning or using any supervised data, consistently outperforms our reproduced implementation of d1. Notably, wd1 surpasses d1 by 43% in test accuracy on the Sudoku task, and achieves up to a 25% improvement on Countdown with maximum length 256. Relative to the base LLaDA model, the performance gain reaches as high as 54% on Sudoku and 42% on Countdown. On math problem-solving benchmarks GSM8K and MATH500, wd1 attains slightly higher accuracy. Nevertheless, the extended method wd1++ obtains significantly better accuracy. In Table˜3 (left), we further compare with concurrent baselines released in recent months. wd1++ outperforms the baselines including strong one MDPO.
Reduced Training Cost. Table˜2 demonstrates that the training cost required by wd1 is substantially lower than that of d1. Unlike d1, wd1 does not require a SFT stage, which alone accounts for approximately two hours of training in d1. wd1 achieves additional speedup during the RL phase, where runtime is measured by averaging over $\mu=8$ inner gradient steps per global step. Notably, the time efficiency gap is expected to widen further under settings with larger maximum sequence lengths and more diffusion steps. This efficiency gain is further supported by a reduced FLOPs and number of function evaluations (NFEs) per step, as wd1 bypasses the need to approximate the likelihood of the old policy. We exclude NFEs associated with sampling, since both methods share identical sampling costs as wd1 removes the reference policy regularization.
In Table˜3 (right), we report the training cost required to obtain the best post-trained models on GSM8K and MATH500, measured in terms of the number of training steps and rollouts. wd1++ requires $10\times$ fewer rollouts to achieve superior performance, clearly demonstrating the efficiency of our method. This rapid convergence arises primarily from the exponential advantage weights applied to the log-likelihood in wd1. In contrast, standard RL methods such as GRPO and PPO weight the log-likelihood (or policy ratio) terms directly by the advantage function.
Table 3: Left: Extended method wd1++ compared to concurrent RL methods to fine-tune LLaDA-8B-Instruct. Methods denoted by “(full)” perform full fine-tuning. Right: Training cost to obtain the best model on GSM8K and MATH500. We count the total number of steps of policy iteration (model weights update), and the number of rollouts used for training (see Table 8 for details on counting).
Model GSM8K MATH500
LLaDA-8B-Instruct 78.2 36.2
+ diffu-GRPO (Zhao et al., 2025) 80.7 39.0
+ d1 (Zhao et al., 2025) 82.0 38.0
+ SDPO (Han et al., 2025) (full) 81.2 -
+ TCR (Wang et al., 2025d) 83.0 41.4
+ MDPO (He et al., 2025) (full) 83.4 43.4
+ wd1 82.3 39.0
+ wd1 (full) 82.7 43.6
+ wd1++ (full) 84.5 44.2
5.2 Ablation Study
Table 4: Ablation on SFT and Negative Samples Weight ($w^{-}$). We conduct wd1 training after SFT (wd1-SFT) and with only $w^{+}$ (namely wd1-P or WLL defined in Equation˜6) [^5] To facilitate efficient ablation studies, we restrict our comparisons to checkpoints saved prior to 5000 steps.. Results show that wd1 performs better without SFT on planning and math tasks. Removing negative sample reinforcement ($w^{-}$) significantly hurts performance, highlighting its importance.
Model / Gen Len Sudoku Countdown GSM8K MATH500
256 512 256 512 256 512 256 512
wd1-P (WLL) 6.69 6.84 13.67 4.69 65.66 78.17 29.40 22.80
wd1-SFT 26.5 24.2 43.4 43.4 80.7 82.0 36.4 39.0
wd1 25.2 24.2 51.2 46.1 80.8 82.3 34.4 39.0
We present an ablation study in Figure 4. Notably, we observe that supervised fine-tuning (SFT) yields only marginal improvements within our approach, with a slight gain in the Sudoku task. This contrasts with d1, where SFT plays a significant role in improving performance. These findings indicate that wd1 can eliminate the need for an SFT phase, thereby simplifying the training pipeline and substantially reducing computational cost. Additionally, we evaluate the impact of removing the negative-weighted term by setting $w^{-}=0$, thus relying solely on the positive advantage weights $w^{+}$. We provide further ablation on the combined method between $w^{+}$ and $w^{-}$ in Table˜9. The results highlight the importance of explicitly penalizing the likelihood of low-advantage completions, thereby reinforcing the role of negative samples, and emphasize the critical balance between the two weights.
Figure 2: Training rewards of wd1 under different combined weights on Sudoku.
We further assess sensitivity to the relative weighting of positive and negative samples. The combined weight (cw) corresponds to $\lambda$ in the mixture $\lambda w^{+}+(1-\lambda)w^{-}$, which scales the log-likelihood term in wd1. Training on negative samples alone (cw=0.0=0.0) yields a pronounced deterioration in performance relative to our default setting (cw=0.5=0.5). The results reinforce our argument that a balanced contribution of positive and negative weights is most effective. In the absence of positive samples, the reinforcement-learning signal collapses and optimisation becomes largely ineffective. A large emphasis on positive samples (cw=0.8=0.8) causes performance to deteriorate more rapidly, highlighting the critical role of negative samples in weighted log-likelihood methods.
6 Related Work
RL for Diffusion-based LLM. RL for discrete diffusion models has been explored through several approaches. One line of work, exemplified by DRAKES (Wang et al., 2024), leverages reward backpropagation along the denoising trajectory. This approach requires computing a critic and propagating gradients through each denoising step, which is computationally intensive and prone to vanishing gradients. Alternatively, methods such as MMaDA (Yang et al., 2025) and d1 (Zhao et al., 2025) adopt direct RL formulations like GRPO, approximating missing diffusion components—such as per-token likelihoods—for policy optimization. Zhu et al. (2025a) applies Direct Preference Optimization (DPO) to fine-tune the LLaDA base model (Nie et al., 2025a), achieving notable gains in reasoning tasks. However, these approaches all depend on likelihood ratios, which can introduce bias and instability due to likelihood approximation errors. In contrast, our method derives a weighted policy optimization approach that eliminates the need for explicit policy ratios. Importantly, similar to prior works, our method directly optimizes the predictive distribution over clean data. A complementary line of research formulates policy optimization in terms of concrete scores (Lou et al., 2024; Meng et al., 2022). SEPO (Zekri and Boullé, 2025), for instance, introduces a policy optimization objective that only depends on concrete score estimation, thereby circumventing likelihood approximation altogether.
RL for AR Models. The connection between GRPO and weighted regression has recently been explored in the context of RL with verifier reward (Mroueh, 2025), where binary rewards simplify policy optimization into likelihood-based objectives. Other closely related approaches are Rejection Sampling Fine-Tuning (RAFT), which maximizes the likelihood of positive-reward samples (Xiong et al., 2025). Extensions of this idea incorporate negative samples to actively penalize the likelihood of negative-reward completions while enhancing that of high-reward ones (Zhu et al., 2025b; Chen et al., 2025). Other works introduce negative penalization through contrastive methods, such as Noise Contrastive Estimation (NCE) (Gutmann and Hyvärinen, 2012; van den Oord et al., 2019; Chen et al., 2024). Beyond binary rewards, preference-based learning has been widely studied using the Bradley–Terry model (Bradley and Terry, 1952; Ouyang et al., 2022; Rafailov et al., 2024; Azar et al., 2023; Ethayarajh et al., 2024; Wang et al., 2023; Hong et al., 2024). In contrast to these approaches, our method accommodates general reward signals and can be interpreted as a form of soft rejection sampling, enabling efficient and stable policy optimization for dLLMs.
RL via Weighted Regression. RL via weighted regression has been explored in earlier works advantage-weighted regression (AWR) (Peng et al., 2019; Peters et al., 2010), and more recently in the context of continuous control with diffusion policies (Ding et al., 2024; Zhang et al., 2025). Weighted likelihood-based approaches have also been proposed for fine-tuning autoregressive (AR) language models using general reward functions (Du et al., 2025; Baheti et al., 2024; Zhu et al., 2023). However, for AR models, where likelihoods are tractable, the necessity of such approaches remains unclear. In contrast, dLLMs suffer from intractable likelihoods, making weighted likelihood formulations particularly advantageous by reducing the number of required likelihood approximations. As such, RL via weighted likelihood provides a natural and efficient fit for optimizing dLLMs. In addition, we demonstrate in ablation study that merely optimizing policy with AWR (wd1-P) is ineffective.
"Ratio-Free" Policy Optimization. If a policy optimization objective requires neither importance sampling nor regularization with respect to a reference model, then the objective is ratio-free. Consequently, on-policy algorithms such as vanilla policy gradient methods (e.g., REINFORCE (Williams, 1992)) and their variants (e.g., RLOO (Kool et al., 2019)) are inherently ratio-free. This property is particularly valuable for dLLMs, where errors in log-likelihood approximation can accumulate and propagate through ratio-based computations. Concurrent work, such as SPG (Wang et al., 2025a), adopts a policy-gradient formulation and develops an objective tailored specifically for diffusion language models. Another on-policy optimization approach, d2 (Wang et al., 2025c), removes both the ratios and the likelihood terms from the RL objective for dLLMs, offering a more fundamental solution. However, our method wd1, similar to AWR (Peng et al., 2019), is inherently an off-policy loss, which is more general.
7 Conclusion
We introduce wd1, a weighted policy optimization method for reasoning with dLLMs. wd1 is designed to minimize reliance on likelihood approximation, thereby mitigating the potentially substantial bias that can arise from approximation errors in policy ratios. Our method is grounded in a weighted log-likelihood objective, derived to approximate the closed-form solution to the reverse-KL-constrained policy optimization. Empirically, we show that wd1, even without supervised fine-tuning, surpasses the existing method d1 by up to 16% in accuracy on reasoning benchmarks, while also delivering notable improvements in computational efficiency during RL training. These results highlight the effectiveness of wd1 and establish it as a more scalable and efficient approach for fine-tuning dLLMs.
8 Ethics and Reproducibility Statement
This work raises no question or concern regarding the Code of Ethics. As for reproducibility of our results, we provide details of implementations in Section˜5, in Experimental Setup and Implementation subsections. Additional details including dataset, reward functions, and hyperparameters are provided in Appendix˜B. All the theoretical results are proved in Appendix˜A.
9 Acknowledgments
Ilija Bogunovic was supported by the EPSRC New Investigator Award EP/X03917X/1; the Engineering and Physical Sciences Research Council EP/S021566/1. Sangwoong Yoon was supported by the Institute of Information & Communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (No. RS-2020-II201336, Artificial Intelligence Graduate School Program (UNIST)), the National Research Foundation of Korea(NRF) grant funded by the Korea government(MSIT) (No. RS-2024-00408003), and the Center for Advanced Computation at Korea Institute for Advanced Study. Xiaohang Tang was supported by the Engineering and Physical Sciences Research Council [grant number EP/T517793/1, EP/W524335/1]. Rares Dolga is supported by EPSRC, grant reference number EP/S021566/1.
The authors would like to thank UIPath and Che Liu (Imperial College London) for providing computing resources that supported our experiments, and Prof. David Barber, Yiming Yang, Xiaoyuan Cheng, and Keyue Jiang from University College London for valuable discussions during the early stages of this work.
Appendix A Proofs and Additional Theory
A.1 Objective Estimation Error due to Likelihood Approximation
In this section, we aim to show that diffu-GRPO amplify the log-likelihood approximation error. Denote the approximator by $\phi$ such that $\|\phi^{\pi_{\theta}}(q,o)-\log\pi_{\theta}(o|q)\|\leq\epsilon$ and $\|\phi^{\pi_{\text{old}}}(q,o)-\log\pi_{\text{old}}(o|q)\|\leq\epsilon^{\prime}$. Then the objective diffu-GRPO in the worst case suffers from exponential error. We discuss the case without ratio clipping and omit the regularization for convenience. Denote $\mathcal{L}_{\text{GRPO}}$ as the ground truth objective without likelihood approximation:
$\|$ $\mathcal{L}_{\text{{diffu}-GRPO}}-\mathcal{L}_{\text{GRPO}}\|$
$=$ $||\mathbb{E}_{q\sim\mathcal{D},\,o_{1:G}\sim\pi_{\text{old}}(\cdot\mid q)}\Bigg[\tfrac{1}{G}\sum_{i=1}^{G}\tfrac{1}{|o_{i}|}\sum_{k=1}^{|o_{i}|}\big(\exp{\phi^{\pi_{\theta}}(o_{i}^{k})}/\exp{\phi^{\pi_{\text{old}}}(o_{i}^{k})}\big)\hat{A}_{i}\big]\Bigg]$
- $\mathbb{E}_{q\sim\mathcal{D},\,o_{1:G}\sim\pi_{\text{old}}(\cdot\mid q)}\Bigg[\tfrac{1}{G}\sum_{i=1}^{G}\tfrac{1}{|o_{i}|}\sum_{k=1}^{|o_{i}|}\big(\pi_{\theta}(o_{i}^{k})/\pi_{\text{old}}(o_{i}^{k})\big)\hat{A}_{i}\Bigg]||$
$=$ $||\mathbb{E}_{q\sim\mathcal{D},\,o_{1:G}\sim\pi_{\text{old}}(\cdot\mid q)}\Bigg[\tfrac{1}{G}\sum_{i=1}^{G}\tfrac{1}{|o_{i}|}\sum_{k=1}^{|o_{i}|}\exp{\big(\phi^{\pi_{\theta}}(o_{i}^{k})-\phi^{\pi_{\text{old}}}(o_{i}^{k})\big)}\hat{A}_{i}\big]\Bigg]$
- $\mathbb{E}_{q\sim\mathcal{D},\,o_{1:G}\sim\pi_{\text{old}}(\cdot\mid q)}\Bigg[\tfrac{1}{G}\sum_{i=1}^{G}\tfrac{1}{|o_{i}|}\sum_{k=1}^{|o_{i}|}\big(\pi_{\theta}(o_{i}^{k})/\pi_{\text{old}}(o_{i}^{k})\big)\hat{A}_{i}\big]\Bigg]||$
$\leq$ $||\mathbb{E}_{q\sim\mathcal{D},\,o_{1:G}\sim\pi_{\text{old}}(\cdot\mid q)}\Bigg[\tfrac{1}{G}\sum_{i=1}^{G}\tfrac{1}{|o_{i}|}\sum_{k=1}^{|o_{i}|}\exp{\big(\log\pi_{\theta}(o_{i}^{k})-\log\pi_{\text{old}}(o_{i}^{k})+(\epsilon+\epsilon^{\prime})\big)}\hat{A}_{i}\big]\Bigg]$
- $\mathbb{E}_{q\sim\mathcal{D},\,o_{1:G}\sim\pi_{\text{old}}(\cdot\mid q)}\Bigg[\tfrac{1}{G}\sum_{i=1}^{G}\tfrac{1}{|o_{i}|}\sum_{k=1}^{|o_{i}|}\big(\pi_{\theta}(o_{i}^{k})/\pi_{\text{old}}(o_{i}^{k})\big)\hat{A}_{i}\big]\Bigg]||$
$=$ $||\mathbb{E}_{q\sim\mathcal{D},\,o_{1:G}\sim\pi_{\text{old}}(\cdot\mid q)}\Bigg[\tfrac{1}{G}\sum_{i=1}^{G}\tfrac{1}{|o_{i}|}\sum_{k=1}^{|o_{i}|}\exp{\big(\epsilon+\epsilon^{\prime}\big)}\hat{A}_{i}\big]\Bigg]\leq C\exp{\big(\epsilon+\epsilon^{\prime}\big)},$ (15)
where $C$ is a constant independent to $\epsilon$ and $\epsilon^{\prime}$. In contrast wd1 has only linear approximation error. Denote the objective computed using approximated log-likelihood as $\mathcal{L}_{\phi}$
$\|\mathcal{L}_{\phi}-\mathcal{L}_{\textit{wd1}}\|=$ $\|\mathbb{E}_{q\sim\mathcal{D},\{o_{i}\}_{i=1}^{G}\sim\pi_{\text{old}}^{\text{ref}}(\cdot|q)}\Big[\sum_{i=1}^{G}\big(-w^{+}(q,o_{i})+w^{-}(q,o_{i})\big)\cdot\big(\phi(q,o_{i})-\log\pi_{\theta}(o_{i}|q)\big)\Big]\|$
$\leq$ $C^{\prime}\epsilon.$ (16)
A.2 Reinforcement Learning
Reinforcement Learning Formulation. We first introduce the reinforcement learning notations and then extend it to the setting of LLM post-training. Denote $\tau$ as a trajectory ($\tau=(s_{0},a_{0},s_{1},\dots)\sim\pi$) sampled following policy $\pi$. Specifically, $s_{0}\sim\mu$, $a_{t}\sim\pi(\cdot|s_{t})$, $s_{t+1}\sim P(\cdot|q_{t},a_{t})$. The objective of Reinforcement Learning aims to find policy $\pi$, which maximizes a discounted total return,
$\eta(\pi)=\mathbb{E}_{\tau\sim\pi}\big[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t},s_{t+1})\big].$
Let the discounted return of a trajectory be $R(\tau)=\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t},s_{t+1})$. The advantage function is defined as $A^{\pi}(s,a)=Q^{\pi}(s,a)-V^{\pi}(s)$ , where $V^{\pi}(s)=\mathbb{E}_{\tau\sim\pi}[R(\tau)|s_{0}=s]$ is state value function, and $Q^{\pi}(s,a)=\mathbb{E}_{\tau\sim\pi}[R(\tau)|s_{0}=s,a_{0}=a]$ is state-action value function. Denote $\rho_{\pi_{\text{old}}}$ as the marginal state distribution. Denote the total variation of two discrete probability distributions $a,b$ by $D_{TV}(a,b):=\frac{1}{2}\sum_{i}|a_{i}-b_{i}|$ and $D_{TV}(a,b)^{2}\leq D_{\text{KL}}(a\parallel b)$ (Pollard, 2000; Schulman et al., 2015). When $a$ and $b$ are conditional probability distribution, denote $D_{TV}^{\text{max}}(a,b)=\max_{q}D_{\text{TV}}(a(\cdot|q)\|b(\cdot|q))$ and $D^{\text{max}}_{\text{KL}}(a\parallel b)=\max_{q}D_{\text{KL}}(a(\cdot|q)\|b(\cdot|q))$.
We then extend RL for LLM post-training. In this paper we only consider the sequence-level reward and loss objective, so we directly replace $s$ with $q$ and $a$ with completion $o$. Then the horizon of the RL for post-training becomes only $1$. The following theorem provides a monotonic (non-decreasing) guarantee of existing prevailing RL methods.
Proposition 1 (Policy Improvement Bound (Kakade and Langford, 2002; Schulman et al., 2015)).
Let surrogate objective $L_{\pi_{\text{old}}}(\pi)=\eta(\pi_{\text{old}})+\mathbb{E}_{s\sim\rho_{\pi_{\text{old}}}(\cdot),\ a\sim\pi(\cdot\mid s)}\big[A^{\pi_{\text{old}}}(s,a)\big]$, and $C={4\max_{s,a,\pi}|A^{\pi}(s,a)|\gamma}/{(1-\gamma)^{2}}$, then $\forall k\in\mathbb{N}$:
$\eta(\pi^{*})\geq L_{\pi_{\text{old}}}(\pi^{*})-CD_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\text{TV}}}^{\max}(\pi_{\text{old}},\pi^{*})^{2}.$
Remark 3.
Based on Proposition˜1, due to $D_{\text{TV}}^{\max}(a||b)^{2}\leq D_{\text{KL}}^{\max}(a||b)$ (Pollard, 2000; Schulman et al., 2015), TRPO and PPO with fixed forward KL regularization have the monotonic improvement guarantees. In other words, $\eta(\pi^{*})\geq L_{\pi_{\text{old}}}(\pi^{*})-CD_{{\text{TV}}}^{\max}(\pi_{\text{old}},\pi^{*})^{2}\geq L_{\pi_{\text{old}}}(\pi^{*})-C\mathbb{E}[D_{\text{KL}}(\pi_{\text{old}}\|\pi^{*})]\geq L_{\pi_{\text{old}}}(\pi_{\text{old}})=\eta(\pi_{\text{old}})$.
Proposition 2.
Minimizing $D_{\text{KL}}(\pi^{*}(\cdot|q)\,\|\,\pi_{\theta}(\cdot|q))$ w.r.t. $\theta$ is equivalent to optimize the following loss objective:
$\mathcal{L}_{\text{WLL}}(\theta)=$ $\mathbb{E}_{q\sim\mathcal{D},o\sim\pi^{\text{ref}}_{\text{old}}(\cdot|q)}\Big[-\exp\big(\psi A^{\pi_{\text{old}}}(q,o)\big)\cdot\log\pi_{\theta}(o_{i}|q)\Big]$ (17)
$\approx$ $-\mathbb{E}_{\{o_{i}\}_{i=1}^{\mathrm{G}}\sim\pi_{\text{old}}^{\text{ref}}(o|q)}\Big[\frac{1}{G}\sum_{i=1}^{\mathrm{G}}\frac{\exp\big(\psi A^{\pi_{\text{old}}}(q,o_{i})\big)}{\sum_{j=1}^{\mathrm{G}}[\exp\big(\psi A^{\pi_{\text{old}}}(q,o_{j})\big)]}\cdot\log\pi_{\theta}(o_{i}|q)\Big].$ (18)
Proof.
To obtain the practical objective in Equation˜18, we first start from the cross-entropy loss, and obtain the following. $\forall q\in\mathcal{D}$:
$D_{\text{KL}}(\pi^{*}(\cdot|q)\,\|\,\pi_{\theta}(\cdot|q))$
$=-\mathbb{E}_{o\sim\pi^{*}(\cdot|q)}\Big[\log\pi_{\theta}(o|q)\Big]$ (19)
$=-\Big[\sum_{o}\pi^{*}(o|q)\cdot\log\pi_{\theta}(o|q)\Big]$ (20)
$=-\Big[\sum_{o}\frac{\pi_{\text{old}}^{\text{ref}}(o|q)\exp\big(\psi A^{\pi_{\text{old}}}(q,o)\big)}{\sum_{o^{\prime}}\pi_{\text{old}}^{\text{ref}}(o^{\prime}|q)[\exp\big(\psi A^{\pi_{\text{old}}}(q,o^{\prime})\big)]}\cdot\log\pi_{\theta}(o|q)\Big]$ (21)
$=-\mathbb{E}_{o\sim\pi_{\text{old}}^{\text{ref}}(o|q)}\Big[\frac{\exp\big(\psi A^{\pi_{\text{old}}}(q,o)\big)}{\mathbb{E}_{o^{\prime}\sim\pi_{\text{old}}^{\text{ref}}}[\exp\big(\psi A^{\pi_{\text{old}}}(q,o^{\prime})\big)]}\cdot\log\pi_{\theta}(o|q)\Big]$ (22)
Since the normalization constant $\mathbb{E}_{o^{\prime}\sim\pi_{\text{old}}^{\text{ref}}}[\exp\big(\psi A^{\pi_{\text{old}}}(q,o^{\prime})\big)]$ is independent to $o$, we can convert the objective to a weighted log-likelihood, and approximate it with samples from the group and weight normalization to obtain:
$\mathcal{L}_{\text{WLL}}(\theta)=$ $-\mathbb{E}_{o\sim\pi_{\text{old}}^{\text{ref}}(o|q)}\Big[{\exp\big(\psi A^{\pi_{\text{old}}}(q,o)\big)}\cdot\log\pi_{\theta}(o|q)\Big]$ (23)
$\approx$ $-\mathbb{E}_{\{o_{i}\}_{i=1}^{\mathrm{G}}\sim\pi_{\text{old}}^{\text{ref}}(o|q)}\Big[\frac{1}{G}\sum_{i=1}^{\mathrm{G}}\frac{\exp\big(\psi A^{\pi_{\text{old}}}(q,o_{i})\big)}{\sum_{j=1}^{\mathrm{G}}[\exp\big(\psi A^{\pi_{\text{old}}}(q,o_{j})\big)]}\cdot\log\pi_{\theta}(o_{i}|q)\Big].$ (24)
We derive Equation˜21 from Equation˜20 by simply using the known form of the optimal policy $\pi^{*}(\cdot|q)\propto\pi_{{\text{old}}}^{\text{ref}}(\cdot|q)\cdot\exp\big(\psi\hat{A}(q,\cdot)\big)$. We derive Equation˜22 from Equation˜21 by using the definition of expectation and from Equation˜22 to Equation˜24 by approximating through $G$ samples $\{o_{i}\}_{i=1}^{\mathrm{G}}\sim\pi_{\text{old}}^{\text{ref}}(o|q)$.
Theorem 2.
Reverse-KL-regularized Policy Optimization defined in the following objective has monotonic improvement guarantees. Specifically, denote regularized objective $\eta^{\prime}(\pi)=\eta(\pi)-\mathbb{E}_{q\in\mathcal{D}}\big[\beta D_{\text{KL}}\big(\pi(\cdot|q)\,\|\,\pi_{{\text{ref}}}(\cdot|q)\big)\big]$ and denote
$M(\pi)=L(\pi)-\mathbb{E}_{q\in\mathcal{D}}\Big[\lambda D_{\text{KL}}(\pi(\cdot|q)\|\pi_{\text{old}}(\cdot|q))+\beta D_{\text{KL}}\Big(\pi(\cdot|q)\,\|\,\pi_{{\text{ref}}}(\cdot|q)\Big)\Big],$ (25)
where $L(\pi)=\eta(\pi_{\text{old}})+\mathbb{E}_{q\sim\mathcal{D},\ o\sim\pi(\cdot\mid q)}\big[A^{\pi_{\text{old}}}(q,o)\big]$. Let $\theta^{*}$ be the solution to the objective $\max_{\theta}M(\pi_{\theta})$:
[…]
then $\eta^{\prime}(\pi^{*})\geq\eta^{\prime}(\pi_{\text{old}})$.
Proof.
Based on Proposition˜1, we have
$\eta^{\prime}(\pi^{*})=$ $\eta(\pi^{*})-\mathbb{E}_{q\in\mathcal{D}}\Big[\beta D_{\text{KL}}\Big(\pi^{*}(\cdot|q)\,\|\,\pi_{{\text{ref}}}(\cdot|q)\Big)\Big]$
$\geq$ $L(\pi^{*})-CD_{\text{TV}}^{\max}(\pi_{\text{old}},\pi^{*})^{2}-\mathbb{E}_{q\in\mathcal{D}}\Big[\beta D_{\text{KL}}\Big(\pi^{*}(\cdot|q)\,\|\,\pi_{{\text{ref}}}(\cdot|q)\Big)\Big]$ (27)
$\geq$ $L(\pi^{*})-CD_{\text{KL}}^{\max}(\pi^{*}\|\pi_{\text{old}})-\mathbb{E}_{q\in\mathcal{D}}\Big[\beta D_{\text{KL}}\Big(\pi^{*}(\cdot|q)\,\|\,\pi_{{\text{ref}}}(\cdot|q)\Big)\Big]$ (28)
$\geq$ $L(\pi^{*})-\mathbb{E}_{q\in\mathcal{D}}\Big[\lambda D_{\text{KL}}(\pi^{*}(\cdot|q)\|\pi_{\text{old}}(\cdot|q))+\beta D_{\text{KL}}\Big(\pi^{*}(\cdot|q)\,\|\,\pi_{{\text{ref}}}(\cdot|q)\Big)\Big]$ (29)
$=$ $M(\pi^{*})$ (30)
$\geq$ $M(\pi_{\text{old}})$ (31)
$=$ $L(\pi_{\text{old}})-\mathbb{E}_{q\in\mathcal{D}}\Big[\lambda D_{\text{KL}}(\pi_{\text{old}}(\cdot|q)\|\pi_{\text{old}}(\cdot|q))+\beta D_{\text{KL}}\Big(\pi_{\text{old}}(\cdot|q)\,\|\,\pi_{{\text{ref}}}(\cdot|q)\Big)\Big]$ (32)
$\geq$ $L(\pi_{\text{old}})-\mathbb{E}_{q\in\mathcal{D}}\Big[\beta D_{\text{KL}}\Big(\pi_{\text{old}}(\cdot|q)\,\|\,\pi_{{\text{ref}}}(\cdot|q)\Big)\Big]$ (33)
$=$ $\eta(\pi_{\text{old}})-\mathbb{E}_{q\in\mathcal{D}}\Big[\beta D_{\text{KL}}\Big(\pi_{\text{old}}(\cdot|q)\,\|\,\pi_{{\text{ref}}}(\cdot|q)\Big)\Big]$ (34)
$=$ $\eta^{\prime}(\pi_{\text{old}})$ (35)
Equation˜27 holds due to Proposition˜1. Equation˜28 holds due to $D_{\text{TV}}^{\max}(p||q)^{2}\leq D_{\text{KL}}^{\max}(p||q)$ (Pollard, 2000). Equation˜29 holds due to the definition of $D^{\text{max}}_{\text{KL}}$. Equation˜30 is according to the definition of $M(\cdot)$. The key inequality Equation˜31 holds since $\pi^{*}$ is the maximizer of function $L(\cdot)$. Equation˜32 holds due to the definition of $M(\cdot)$. Equation˜33 holds since $D_{\text{KL}}(\pi_{\text{old}}(\cdot|q)\|\pi_{\text{old}}(\cdot|q))=0$. Equation˜34 holds since $L(\pi_{\text{old}})=\eta(\pi_{\text{old}})+\mathbb{E}_{q\sim\mathcal{D},\ o\sim\pi_{\text{old}}(\cdot\mid q)}\big[A^{\pi_{\text{old}}}(q,o)\big]=\eta(\pi_{\text{old}})$. Equation˜35 is from the definition of $\eta^{\prime}$. ∎
A.3 Masked Discrete Diffusion
In this section, we show how our objective learns a distribution for which all marginals at time $t$ satisfy intermediate energy guidance as per Lu et al. (2023).
Definition 2.
The absorbing transition kernel is defined as $Q_{t}=\sigma(t)Q^{\text{absorb}}$, where
$Q^{\text{absorb}}=\begin{bmatrix}-1&0&\cdots&0&1\\ 0&-1&\cdots&0&1\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ 0&0&\cdots&-1&1\\ 0&0&\cdots&0&0\end{bmatrix}$.
Definition 3 (Concrete Score).
Denote $x_{t}=(x_{t}^{1},\cdots,x_{t}^{d})$ and $\hat{x}_{t}$ is identical to $x_{t}$ except the $i$-th token is unmasked (i.e. $x_{t}^{i}=[M]$ and $\hat{x}_{t}^{i}\neq[M]$). Concrete score is defined as the marginal probability ratio between $\hat{x}_{t}$ and $x_{t}$:
$s(x_{t},t)\overset{\text{def}}{=}\frac{p(x_{t}^{1},\cdots,\hat{x}_{t}^{i},\cdots,x_{t}^{d})}{p(x_{t}^{1},\cdots,x_{t}^{i},\cdots,x_{t}^{d})}.$ (36)
Proposition 3 (Marginal Distribution (Ou et al., 2025)).
Denote $\{x_{t}\}$ as a continuous time Markov chain with transition matrix $\mathbf{Q}_{t}=\sigma(t)\mathbf{Q}^{\text{absorb}}$. Assume $d_{1}$ tokens in $x_{t}=(x_{t}^{1},\cdots,x_{t}^{d})$ are masked tokens $\mathbf{M}]$, and $d_{2}=d-d_{1}$ tokens are unmasked, the marginal distribution $p_{t}(x_{t})$ satisfies
$p_{t}(x_{t})=\big[1-e^{-\bar{\sigma}(t)}\big]^{d_{1}}\big[e^{-\bar{\sigma}(t)}\big]^{d_{2}}p_{0}(x_{t}^{\text{UM}}),$ (37)
where $\bar{\sigma}(t)=\int_{0}^{t}\sigma(s)ds$, and $x_{t}^{\text{UM}}$ is the set of unmasked tokens in $x_{t}$.
The following theorem provides the foundation of directly modeling the clean data distribution.
Proposition 4 (Analytic Concrete Score  (Ou et al., 2025)).
Denote $x_{t}=(x_{t}^{1},\cdots,x_{t}^{d})$ and $\hat{x}_{t}$ is identical to $x_{t}$ except the $i$-th token is unmasked (i.e. $x_{t}^{i}=[M]$ and $\hat{x}_{t}^{i}\neq[M]$). Then the concrete score at time $t$ can be expressed by the conditional probability of predicting this unmasked token.
$\frac{p_{t}(x_{t}^{1}\ldots\hat{x}_{t}^{i}\ldots x_{t}^{d})}{p_{t}(x_{t}^{1}\ldots x_{t}^{i}\ldots x_{t}^{d})}=\frac{e^{-\bar{\sigma}(t)}}{1-e^{-\bar{\sigma}(t)}}p_{0}(\hat{x}_{t}^{i}\mid x_{t}^{\text{UM}})$
Lemma (1).
The marginal probability distribution of the masked responses ($x_{t}$) in the diffusion process satisfies $p^{*}_{t}(x_{t})=p^{\prime}_{t}(x_{t})\cdot\exp\big(A_{t}(x_{t})\big)/Z_{t}$, which induces an energy-guided discrete diffusion:
$p^{*}_{0|t}(x_{0}|x_{t})\propto p^{\prime}_{0|t}(x_{0}|x_{t})\cdot\exp(A(x_{0})-A_{t}(x_{t})),$ (38)
where intermediate energy function is defined as $A_{t}(x_{0})=\log\mathbb{E}_{x_{0}\sim p^{\prime}_{0|t}(\cdot|x_{t})}[\exp\big(A(x_{0})\big)]$ for $t>0$, and $A_{0}(x_{0})=A(x_{0})$, $A(\cdot)$ is advantage function, $Z_{t}$ is the normalization constant.
Proof.
The theorem and proof mainly extend from theory developed in continuous setting (Lu et al., 2023). According to the marginal likelihood of clean data distribution $p^{*}_{0}(x_{0})=p^{\prime}_{0}(x_{0})\frac{e^{A(x_{0})}}{Z}$, and identical forward process, we can rewrite the marginal likelihood of masked data:
$p^{*}_{t}(x_{t})$ $=\int p^{*}_{t|0}(x_{t}|x_{0})p^{*}_{0}(x_{0})\,\mathrm{d}x_{0}=\int p^{*}_{t|0}(x_{t}|x_{0})p^{\prime}_{0}(x_{0})\frac{e^{\psi A(x_{0})}}{Z}\,\mathrm{d}x_{0}$
$=\int p^{\prime}_{t|0}(x_{t}|x_{0})p^{\prime}_{0}(x_{0})\frac{e^{\psi A(x_{0})}}{Z}\,\mathrm{d}x_{0}.$
Applying Bayesian rule we know that $p^{\prime}_{t|0}(x_{t}|x_{0})p^{\prime}_{0}(x_{0})=p^{\prime}_{0|t}(x_{0}|x_{t})p^{\prime}_{t}(x_{t})$, hence we can further rewrite
$p^{*}_{t}(x_{t})$ $=\int p^{\prime}_{t|0}(x_{t}|x_{0})p^{\prime}_{0}(x_{0})\frac{e^{\psi A(x_{0})}}{Z}\,\mathrm{d}x_{0}=p^{\prime}_{t}(x_{t})\int p^{\prime}_{0|t}(x_{0}|x_{t})\frac{e^{\psi A(x_{0})}}{Z}\,\mathrm{d}x_{0}$
$=\frac{p^{\prime}_{t}(x_{t})\,\mathbb{E}_{p^{\prime}_{0|t}(x_{0}|x_{t})}\!\left[e^{\psi A(x_{0})}\right]}{Z}=\frac{p^{\prime}_{t}(x_{t})\,e^{\psi A_{t}(x_{t})}}{Z_{t}}$
Therefore, the marginal likelihood of masked sequence satisfies: $p^{*}_{t}(x_{t})=p^{\prime}_{t}(x_{t})\cdot\exp\big(A_{t}(x_{t})\big)/Z_{t}$. Since $p^{*}_{t|0}=p^{\prime}_{t|0}$, based on the marginal likelihood of clean data distribution satisfies $p^{*}_{0}(x_{0})=p^{\prime}_{0}(x_{0})\frac{e^{A(x_{0})}}{Z}$, we can further applying Bayesian rule to obtain the energy-guided discrete diffusion model:
$p^{*}_{0|t}(x_{0}|x_{t})=$ $\frac{p^{*}_{t|0}(x_{t}|x_{0})p^{*}_{0}(x_{0})}{p^{*}_{t}(x_{t})}$ (39)
$=\frac{p^{\prime}_{t|0}(x_{t}|x_{0})p^{\prime}_{0}(x_{0})\frac{e^{A(x_{0})}}{Z}}{p^{*}_{t}(x_{t})}$ (40)
$=\frac{p^{\prime}_{t|0}(x_{t}|x_{0})p^{\prime}_{0}(x_{0})\frac{e^{A(x_{0})}}{Z}}{\frac{p^{\prime}_{t}(x_{t})\,e^{\psi A_{t}(x_{t})}}{Z_{t}}}$ (41)
$\propto p^{\prime}_{0|t}(x_{0}|x_{t})\cdot\exp(A(x_{0})-A_{t}(x_{t})),$ (42)
Lemma 2.
According to Definition 1, due to the identical forward process
$p^{*}_{t|0}(x_{t}|x_{0})=p^{\prime}_{t|0}(x_{t}|x_{0})=p^{\text{ref}}_{t|0}(x_{t}|x_{0}),$ (43)
based on Lemma˜1 and Proposition˜3, we have the marginal probability of the unmasked tokens satisfies that for all step $t$,
$p^{*}_{0}(x_{t}^{\text{UM}}|q)=$ $p_{0}(x_{t}^{\text{UM}}|q)^{\lambda}\cdot p^{{\text{ref}}}_{0}(x_{t}^{\text{UM}}|q)^{\beta}\cdot\mathbb{E}_{p^{\prime}_{0}(x_{0}|x_{t})}[\exp\big(A(q,x_{0})\big)]/Z,$ (44)
where $Z$ is the normalization constant.
Proof.
According to the identical forward distribution of three diffusion process (new, old, and reference), based on Equation˜37, we have $\forall t$:
$p_{t}(x_{t}|q)$ $=\big[1-e^{-\bar{\sigma}(t)}\big]^{d_{1}}\big[e^{-\bar{\sigma}(t)}\big]^{d_{2}}p_{0}(x_{t}^{\text{UM}}|q)$ (45)
$p^{*}_{t}(x_{t}|q)$ $=\big[1-e^{-\bar{\sigma}(t)}\big]^{d_{1}}\big[e^{-\bar{\sigma}(t)}\big]^{d_{2}}p^{*}_{0}(x_{t}^{\text{UM}}|q)$ (46)
$p^{{\text{ref}}}_{t}(x_{t}|q)$ $=\big[1-e^{-\bar{\sigma}(t)}\big]^{d_{1}}\big[e^{-\bar{\sigma}(t)}\big]^{d_{2}}p^{{\text{ref}}}_{0}(x_{t}^{\text{UM}}|q)$ (47)
Then rewrite Equation˜46 in the residual energy-based form defined in Equation˜38, we have
$\big[1-e^{-\bar{\sigma}(t)}\big]^{d_{1}}\big[e^{-\bar{\sigma}(t)}\big]^{d_{2}}p^{*}_{0}(x_{t}^{\text{UM}}|q)=p^{*}_{t}(x_{t}|q)=$ $p^{\prime}_{t}(x_{t}|q)\cdot\exp\big(A_{t}(q,x_{t})\big)/Z.$ (48)
By plugging $p^{\prime}_{t}(x_{t}|q)=p_{t}(x_{t}|q)^{\lambda}\cdot p^{{\text{ref}}}_{t}(x_{t}|q)^{\beta}$ and Equation˜45 and Equation˜47 into Equation˜48, we have that the clean data distribution of the unmask tokens at diffusion time $t$ satisfies:
$p^{*}_{0}(x_{t}^{\text{UM}}|q)=$ $p_{0}(x_{t}^{\text{UM}}|q)^{\lambda}\cdot p^{{\text{ref}}}_{0}(x_{t}^{\text{UM}}|q)^{\beta}\cdot\exp\big(A_{t}(q,x_{t})\big)/Z$ (49)
$=$ $p_{0}(x_{t}^{\text{UM}}|q)^{\lambda}\cdot p^{{\text{ref}}}_{0}(x_{t}^{\text{UM}}|q)^{\beta}\cdot\mathbb{E}_{p^{\prime}_{0}(x_{0}|x_{t})}[\exp\big(A(q,x_{0})\big)]/Z.$ (50)
Proposition 5.
The marginal likelihood of the target diffusion model $p^{*}$ satisfies Equation˜38. Consequently, the concrete score of the target diffusion model, denoted by $s^{*}$, can be expressed by the score of the mixture diffusion $p^{\prime}$ and the posterior mean of the advantage:
$s^{*}(x_{t},t)=$ $s^{\prime}(x_{t},t)\cdot\frac{\mathbb{E}_{p^{\prime}_{0}(x_{0}|\hat{x}_{t})}[\exp\big(A(x_{0})\big)]/\hat{Z}}{\mathbb{E}_{p^{\prime}_{0}(x_{0}|x_{t})}[\exp\big(A(x_{0})\big)]/Z},$ (51)
and equivalently
${p_{0}(\hat{x}_{t}^{i}|x_{t}^{\text{UM}},q)^{\lambda}}\cdot{p^{\text{ref}}_{0}(\hat{x}_{t}^{i}|x_{t}^{\text{UM}},q)^{\beta}}\cdot\frac{\mathbb{E}_{p^{\prime}_{0}(x_{0}|\hat{x}_{t})}[\exp\big(A(q,x_{0})\big)]/\hat{Z}}{\mathbb{E}_{p^{\prime}_{0}(x_{0}|x_{t})}[\exp\big(A(q,x_{0})\big)]/Z}.$ (52)
Proof.
According to Lemma 2
$\frac{p^{*}_{0}(x_{t}^{\text{UM}},\hat{x}_{t}^{i}|q)}{p^{*}_{0}(x_{t}^{\text{UM}}|q)}=$ $\frac{p_{0}(x_{t}^{\text{UM}},\hat{x}_{t}^{i}|q)^{\lambda}}{p_{0}(x_{t}^{\text{UM}}|q)^{\lambda}}\cdot\frac{p^{\text{ref}}_{0}(x_{t}^{\text{UM}},\hat{x}_{t}^{i}|q)^{\beta}}{p^{\text{ref}}_{0}(x_{t}^{\text{UM}}|q)^{\beta}}\cdot\frac{\mathbb{E}_{p^{\prime}_{0}(x_{0}|\hat{x}_{t})}[\exp\big(A(q,x_{0})\big)]/\hat{Z}}{\mathbb{E}_{p^{\prime}_{0}(x_{0}|x_{t})}[\exp\big(A(q,x_{0})\big)]/Z}$ (53)
${p^{*}_{0}(\hat{x}_{t}^{i}|x_{t}^{\text{UM}},q)}=$ ${p_{0}(\hat{x}_{t}^{i}|x_{t}^{\text{UM}},q)^{\lambda}}\cdot{p^{\text{ref}}_{0}(\hat{x}_{t}^{i}|x_{t}^{\text{UM}},q)^{\beta}}\cdot\frac{\mathbb{E}_{p^{\prime}_{0}(x_{0}|\hat{x}_{t})}[\exp\big(A(q,x_{0})\big)]/\hat{Z}}{\mathbb{E}_{p^{\prime}_{0}(x_{0}|x_{t})}[\exp\big(A(q,x_{0})\big)]/Z}.$ (54)
Both sides in Equation˜52 multiply $C(t)=\tfrac{e^{-\cdot\bar{\sigma(t)}}}{1-e^{\bar{\sigma(t)}}}$ and based on the analytic form of concrete score introduced in Proposition˜4, $C(t)\cdot{p_{0}(\hat{x}_{t}^{i}|x_{t}^{\text{UM}},q)}=s(x_{t}^{i},t)$. Thus, we have
$C(t)\cdot{p^{*}_{0}(\hat{x}_{t}^{i}|x_{t}^{\text{UM}},q)}$ $=C(t)\cdot{p_{0}(\hat{x}_{t}^{i}|x_{t}^{\text{UM}},q)^{\lambda}}\cdot{p^{\text{ref}}_{0}(\hat{x}_{t}^{i}|x_{t}^{\text{UM}},q)^{\beta}}\cdot\frac{\mathbb{E}_{p^{\prime}_{0}(x_{0}|\hat{x}_{t})}[\exp\big(A(q,x_{0})\big)]/\hat{Z}}{\mathbb{E}_{p^{\prime}_{0}(x_{0}|x_{t})}[\exp\big(A(q,x_{0})\big)]/Z}$ (55)
$s^{*}(x_{t},t)$ $=s^{\prime}(x_{t},t)\cdot\frac{\mathbb{E}_{p^{\prime}_{0}(x_{0}|\hat{x}_{t})}[\exp\big(A(q,x_{0})\big)]/\hat{Z}}{\mathbb{E}_{p^{\prime}_{0}(x_{0}|x_{t})}[\exp\big(A(q,x_{0})\big)]/Z}.$ (56)
Lemma 3.
The normalization constant $Z=\sum_{x_{t}}p^{\prime}(x_{t}|q)\cdot\mathbb{E}_{x_{0}|x_{t}}[\exp A(q,x_{0})]$ is independent to the masked response $x_{t}$. In other words, $Z=\hat{Z}:=\sum_{\hat{x}_{t}}p^{\prime}(\hat{x}_{t}|q)\cdot\mathbb{E}_{x_{0}|\hat{x}_{t}}[\exp A(q,x_{0})]$ for any $\hat{x}_{t}\neq x_{t}$.
Proof.
$Z=$ $\sum_{x_{t}}p^{\prime}(x_{t}|q)\cdot\mathbb{E}_{x_{0}|x_{t}}[\exp A(q,x_{0})]=\sum_{x_{t}}p^{\prime}(x_{t}|q)\cdot\sum_{x_{0}}p^{\prime}(x_{0}|x_{t})\cdot\exp A(q,x_{0})$ (57)
$=$ $\sum_{x_{t}}\sum_{x_{0}}p^{\prime}(x_{0},x_{t}|q)\cdot\exp A(q,x_{0})=\sum_{x_{0}}\sum_{x_{t}}p^{\prime}(x_{0},x_{t}|q)\cdot\exp A(q,x_{0})$ (58)
$=$ $\sum_{x_{0}}p^{\prime}(x_{0}|q)\cdot\exp A(q,x_{0})$ (59)
Thus $Z$ becomes independent to $x_{t}$, leading to that
$Z=\hat{Z}:=\sum_{\hat{x}_{t}}p^{\prime}(\hat{x}_{t}|q)\cdot\mathbb{E}_{x_{0}|\hat{x}_{t}}[\exp A(q,x_{0})]$ (60)
Theorem (1).
The score model $s_{\theta}=s^{*}$ defined in Equation˜52 is satisfied when the following loss objective is minimized. This objective is in a form of advantage-weighted Denoising Concrete Score Matching (D-CSM), which we call AW-D-CSM:
$\mathcal{L}_{\text{AW-D-CSM}}=$ $\mathbb{E}_{p^{\prime}_{0}(x_{0})}[\underbrace{\exp\big(A(q,x_{0})\big)}_{\text{Advantage Weight}}\cdot\underbrace{\mathbb{E}_{t\sim[0,T],p^{\prime}_{t|0}(x_{t}|x_{0})}[\|{s_{\theta}(x_{t},t)}-\frac{p^{\prime}_{0}(\hat{x}_{t}|x_{0})}{p^{\prime}_{0}(x_{t}|x_{0})}\|_{2}^{2}]}_{\mathcal{L}_{\text{D-CSM}}(x_{0})}].$ (61)
Proof.
Denote $s_{\theta}(x_{t},t)=\frac{e^{-\bar{\sigma}(t)}}{1-e^{-\bar{\sigma}(t)}}p_{\theta}(\hat{x}_{t}^{i}\mid x_{t}^{\text{UM}})$ is the concrete score model induced by $p_{\theta}$. According to Lemma˜3, $\hat{Z}=Z$. Then according to Proposition 5, Equation˜52 is equivalent to
${p^{*}_{0}(\hat{x}_{t}^{i}|x_{t}^{\text{UM}},q)}\cdot\mathbb{E}_{p^{\prime}_{0}(x_{0}|x_{t})}[\exp\big(A(q,x_{0})\big)]$
$=$ ${p_{0}(\hat{x}_{t}^{i}|x_{t}^{\text{UM}},q)^{\lambda}}\cdot{p^{\text{ref}}_{0}(\hat{x}_{t}^{i}|x_{t}^{\text{UM}},q)^{\beta}}\cdot\mathbb{E}_{p^{\prime}_{0}(x_{0}|\hat{x}_{t})}[\exp\big(A(q,x_{0})\big)]$ (62)
We aim to update $p_{\theta}(\hat{x}_{t}^{i}|x_{t}^{\text{UM}},q)\rightarrow p_{0}^{*}(\hat{x}_{t}^{i}|x_{t}^{\text{UM}},q)$ to satisfy Equation˜62, thus we can construct a loss function objective by replacing $p^{*}$ with $p_{\theta}$ and construct a $L^{2}$ norm loss
$\mathbb{E}_{p^{\prime}_{0}(x_{0}|x_{t})}[\exp\big(A(q,x_{0})\big)\cdot\|{p_{\theta}(\hat{x}_{t}^{i}|x_{t}^{\text{UM}},q)}-\frac{p^{\prime}_{0}(x_{0}|\hat{x}_{t})}{p^{\prime}_{0}(x_{0}|x_{t})}\cdot{p_{0}(\hat{x}_{t}^{i}|x_{t}^{\text{UM}},q)^{\lambda}}{p^{\text{ref}}_{0}(\hat{x}_{t}^{i}|x_{t}^{\text{UM}},q)^{\beta}}\|_{2}^{2}]$ (63)
$=$ $\mathbb{E}_{p^{\prime}_{0}(x_{0}|x_{t})}[\exp\big(A(q,x_{0})\big)\cdot\|{p_{\theta}(\hat{x}_{t}^{i}|x_{t}^{\text{UM}},q)}-\frac{p^{\prime}_{0}(x_{0}|\hat{x}_{t})}{p^{\prime}_{0}(x_{0}|x_{t})}\cdot{p^{\prime}_{0}(\hat{x}_{t}^{i}|x_{t}^{\text{UM}},q)}\|_{2}^{2}]$ (64)
$=$ $\mathbb{E}_{p^{\prime}_{0}(x_{0}|x_{t})}[\exp\big(A(q,x_{0})\big)\cdot\|{p_{\theta}(\hat{x}_{t}^{i}|x_{t}^{\text{UM}},q)}-\frac{p^{\prime}_{0}(\hat{x}_{t}|x_{0})p^{\prime}_{t}(x_{t})}{p^{\prime}_{0}(x_{t}|x_{0})p^{\prime}_{t}(\hat{x}_{t})}\cdot{p^{\prime}_{0}(\hat{x}_{t}^{i}|x_{t}^{\text{UM}},q)}\|_{2}^{2}]$ (65)
$=$ $\mathbb{E}_{p^{\prime}_{0}(x_{0}|x_{t})}[\exp\big(A(q,x_{0})\big)\cdot\|{s_{\theta}(x_{t},t)}-\frac{p^{\prime}_{0}(\hat{x}_{t}|x_{0})}{p^{\prime}_{0}(x_{t}|x_{0})}\|_{2}^{2}].$ (66)
Appendix B Additional Experiment Setup Details
B.1 Dataset, Training and Evaluation Protocol
As for wd1 and d1, we reproduce d1 by running the official code [^6] https://github.com/dllm-reasoning/d1 without and change, and train our method wd1 evaluated for accuracy of the test datasets at steps 1000, 2500, 5000, 7500 in both GSM8k and MATH; at steps 1000, 2500, 4000, 5000, 12500 in Sudoku; and at 1000, 2500, 4000 in Countdown. We evaluate less checkpoints compared to d1. On the GSM8K, we train models on the train split [^7] https://huggingface.co/datasets/openai/gsm8k and evaluate on the test split. On Countdown, we train on the 3-number subset of the dataset [^8] https://huggingface.co/datasets/Jiayi-Pan/Countdown-Tasks-3to4 from TinyZero (Pan et al., 2025), and evaluate on 256 synthetic 3-number questions provided by Zhao et al. (2025). On Sudoku we use the 4×4 dataset [^9] https://github.com/Black-Phoenix/4x4-Sudoku-Dataset generated by Arel (2025). We train on 1M unique puzzles and evaluate on 256 synthetic ones provided by Zhao et al. (2025). On MATH500, we train models on the train split [^10] https://huggingface.co/datasets/ankner/math-500.
To train wd1++ for evaluating on MATH500, we use dataset provided by (He et al., 2025), which is subsampled from OpenR1 dataset Face (2025). To evaluate on GSM8k, we leverage its train split to conduct wd1++ training. Notably, we leverage a more effective system prompt and Math-Verify (Kydlíček, ) to parse the answers for full-parameter fine-tuning of wd1, wd1++ and MDPO.
B.2 Reward Function
To train wd1 and reproduce d1, we use the reward function defined in (Zhao et al., 2025). For completion, we provide the details as following.
[…]
Countdown. We include three cases: +1.0 if the expression reaches the target using the exact numbers. +0.1 if numbers are correct but target is missed. 0 otherwise.
Sudoku. The reward is the fraction of correctly filled empty cells, focusing on solving rather than copying.
[…]
To train wd1++, we leverage Math-Verify (Kydlíček, ), constructing a simple verifier reward function to evalaute on GSM8K and MATH500.
B.3 Sampling from Geometric Mixture
Although the sampling strategy eliminates the need to approximate the reference policy’s likelihood, it incurs computational overhead, as generating a full completion requires multiple forward passes through the dLLM—compared to a single pass for likelihood estimation. An alternative is to sample from $\pi_{\text{old}}$ and shift the advantage to $\hat{A}_{i}=A^{\pi_{\text{old}}}(q,o_{i})+\beta\log\hat{\pi_{\text{ref}}}/(\lambda+\beta)$, which reintroduces the need for reference policy likelihood approximation. However, policy ratio has been removed, and the reference model can be reused when conducting multiple gradient updates with the same batch of rollouts (off-policy). The increased computational burden is slight.
B.4 Hyperparameters
We provide the hyperparameters of SFT in Table 5 and for wd1 in Table 6.
bacth_size max_length learning_rate grad_accum_steps
Value 1 4096 1e-5 4
Table 5: Hyperparameters of SFT in d1 reproduction.
Parameter wd1 d1
Model and Precision
use_peft true true
torch_dtype bfloat16 bfloat16
load_in_4bit true true
attn_implementation flash_attention_2 flash_attention_2
lora_r 128 128
lora_alpha 64 64
lora_dropout 0.05 0.05
peft_task_type CAUSAL_LM CAUSAL_LM
Training Configuration
seed 42 42
bf16 true true
sync_ref_model True True
ref_model_sync_steps 64 64
adam_beta1 0.9 0.9
adam_beta2 0.99 0.99
weight_decay 0.1 0.1
$\psi$ (Equation˜9) 1.0
max_grad_norm 0.2 0.2
warmup_ratio 0.0001 0.0001
learning_rate 3e-6 3e-6
lr_scheduler_type constant_with_warmup constant_with_warmup
Batching and Evaluation
per_device_train_batch_size 6 6
per_device_eval_batch_size 1 1
gradient_accumulation_steps 2 2
RL
num_generations 6 6
max_completion_length 256 256
max_prompt_length 200 200
block_length 32 32
diffusion_steps 128 128
generation_batch_size 6 6
remasking low_confidence low_confidence
random_masking True True
p_mask_prompt 0.15 0.15
beta 0.00 0.04
epsilon 0.5
num_iterations 12 12
Table 6: Comparison of hyperparameters between wd1 and d1.
B.5 Training Cost Estimation
For the runtime measurements reported in Table˜2, we set $\mu=8$ and train for a total of 6 global steps, corresponding to 48 gradient update steps. We use a batch size of 4 and the rest of the hyperparameters are the same as in Table˜6. To estimate the number of function evaluations (NFEs) involved in computing likelihood approximations, we count only the forward passes, as the number of backward passes remains consistent across methods. The additional NFEs observed in the d1 model arise from evaluating the likelihood under both the old and reference models, which are used for regularization. These extra evaluations are required only when new samples are drawn, as their outputs can be cached and reused across all gradient updates for $\mu$. We additionally report the number of floating-point operations (FLOPs) per global training step, measured using the Flops Profiler from Rasley et al. (2020).
B.6 Computing Resources
For both wd1 and d1, RL training is conducted on four NVIDIA A100 GPUs (80GB), and SFT is performed on four A6000 GPUs (48GB). For wd1++ and MDPO, RL training is conducted on $\timesA8$00 (80GB).
Appendix C Additional Experiments
We additionally report results for comparison to the results of the baseline d1 reported in the paper (Zhao et al., 2025). As shown in  Table˜7, our method wd1 evaluated and selected from less checkpoints, can outperform d1 with a large margin in Sudoku and Countdown, achieving comparable performance in math problem-solving tasks.
C.1 Summary of wd1  Results
Table 7: Test accuracy across different tasks. Our method demonstrates higher accuracy, especially significant in Sudoku and Countdown. The shaded area indicates where our method outperforms.
Model Sudoku Countdown GSM8K MATH500
256 512 256 512 256 512 256 512
LLaDA-8B-Instruct 6.7 5.5 19.5 16.0 76.7 78.2 32.4 36.2
+ diffu-GRPO (reported) 12.9 11.0 31.3 37.1 79.8 81.9 37.2 39.2
+ diffu-GRPO (reproduced) 16.1 11.7 27.0 34.0 80.7 79.1 34.4 39.0
d1 (reported) 16.7 9.5 32.0 42.2 81.1 82.1 38.6 40.2
d1 (reproduced) 17.6 16.2 25.8 35.2 78.2 82.0 34.4 38.0
wd1 76.4 62.8 51.2 46.1 80.8 82.3 34.4 39.0
Table 8: Training cost. The training steps to obtained the best post-trained model of three methods are 20, 150, and 7500. To compute the total rollouts, we need to compute the average rollouts in a single training step. Gradient steps per rollout batch represents the number of gradient descent conducted with a single batch of rollouts. In other words, 1 represents it is a pure on-policy RL training, and for any value >12>12, off-policy RL is executed. Total Batch Size is computed by multiplying per-device batch size, gradient accumulation and the number of gpus. Therefore, the average number of rollouts used for single step gradient descent should be computed by total batch size divided by gradient steps per rollout batch.
Hyperparameter wd1++ MDPO d1
Training step of the best checkpoint 20 150 7500
Training Steps per Rollout batch 1 1 12
Per-Device Batch Size 4 1 6
Gradient Accumulation 2 16 2
GPUs used for training 8 8 4
Total Batch Size 64 128 48
Avg. Rollouts per Step 64 128 4
Total Rollouts 1280 19200 30000
We additionally provide reward dynamics in comparison to wd1-SFT in training. In Sudoku and Countdown, directly training with wd1 without SFT shows significantly more efficient and stable learning process. In GSM8k and MATH500, the difference is negligible.
Figure 3: Reward Dynamics. wd1 without SFT demonstrates better rewards in Sudoku and Countdown.
C.2 Additional Ablation Study
Table 9: Ablation on weight $\lambda$ to combine positive $w^{+}$ and negative weights $w^{-}$ in wd1 on Sudoku. Specifically, the final weight assigned to log-likelihood is computed as $\lambda w^{+}+(1-\lambda)w^{-}$.
Combined Weight Accuracy Effective Tokens
0.5 25.63% 326.97
0.4 11.77% 240.04
0.6 14.11% 220.13
We provide additional ablation study on the combined weight to confirm our analysis that the positive and negative samples terms in the loss function should be assigned equal proportion, due to the side case of a batch of all-negative generated responses (see the paragraph below Equation˜9). Assigning equal proportions to positive and negative weights is not arbitrary but rather the most robust design. This can be understood through two critical failure modes that arise from imbalanced proportions:
  • When positive weight has larger proportion: In scenarios where all sampled completions have uniformly low rewards, a larger proportion of positive weights would paradoxically increase the log-likelihood of negative samples during wd1 optimization, which is clearly undesirable and contradicts the learning objective.
  • When negative weight has larger proportion: Conversely, when all generated completions achieve uniformly high rewards, an insufficient proportion of positive weights would result in unlearning high-quality samples.
To empirically validate this analysis, we conducted experiments on the Sudoku dataset with varying mixing proportions. The results, presented in the table below, confirm our theoretical predictions.
Figure 4: Left: Ablation study on $\psi$ in the weights (Equation˜9) on GSM8K. Right: wd1 training on MATH with a random seed different from the seed used in our main experiments. The abrupt decrease of the rewards in the early training (see Figure˜3 MATH500) disappears.
In all benchmark evaluations, we fix the hyperparameter $\psi=1$, which controls the scale of the exponential weighting in wd1. To validate this choice, we provide an ablation study on the coefficient $\psi=1$ in the exponential weight of wd1 (Equation 9) below. Larger values $\psi$ leads to more extreme weight assigned to the samples. According to Figure˜4, the training of applying different $\psi$ converges to similar rewards if $\psi$ is small. Overly large value (e.g. 10) can cause performance drop, implying that extreme weight assignment is detrimental.
C.3 Coding Benchmarks
We conduct 200 steps of wd1 training on AceCode-87K (Zeng et al., 2025) following the implementation of Open-R1 (Hugging Face, 2025). Our method achieves consistent improvements over the base model.
Table 10: Comparative performance of wd1 improvements compared to base model LLaDA-8B-Instruct. We present the results of wd1 fine-tuned with AceCode dataset (Zeng et al., 2025).
Task Gen Length Steps Block Size wd1 LLaDA
HumanEval 256 128 32 34.76 (+3.66) 31.10
HumanEval 256 256 32 39.02 (+1.82) 37.20
HumanEval 512 512 32 36.59 (+0.61) 35.98
MBPP 128 128 32 39.2 (+2.40) 36.8
MBPP 256 256 32 36.6 (+1.00) 35.6
MBPP 512 512 32 36.8 (+0.40) 36.4
C.4 Training Dynamics
Figure˜3 presents the reward dynamics over gradient steps during training. wd1 exhibits a notably faster reward increase compared to d1, highlighting its superior sample efficiency–effectively leveraging the reward signal to accelerate policy optimization. In addition, Figure˜5 shows the average length of generated completions during training. On math reasoning benchmarks such as GSM8K and MATH500, wd1 converges to shorter output sequences than d1, suggesting improved token efficiency while maintaining or improving performance.
Figure 5: Completion lengths dynamics of wd1 and d1. In math problem-solving tasks (GSM8K and MATH500), our method demonstrates smaller completion lengths and better token efficiency.
Appendix D Limitations
Similar to other RL-based approaches, wd1 may lose effectiveness when all generations within a sampled group receive identical rewards. This situation can occur under several conditions—for example, when the training dataset is either too simple or too challenging for the base model. Nonetheless, such cases can be mitigated through careful reward design and the incorporation of curriculum learning strategies.
An additional limitation of this work is that the current wd1 framework is restricted to text-based reasoning. Extending it to multimodal reasoning or unified diffusion-based models (e.g., (Yang et al., 2025)) represents a valuable direction for future research.
A final limitation concerns the likelihood approximation used in wd1. Our approach relies on the d1-based approximation, which is computationally efficient but introduces bias. Although some prior works employ ELBO-based estimators (e.g., DCE), they require additional computational overhead (Zhao et al., 2025) often exhibit high variance, as demonstrated in Figure 1. This trade-off highlights an important area for further exploration.
D.1 Additional Analysis on Unlearning
We provide extended demonstrations for Remark Remark˜2, focusing specifically on the theoretical insights underlying the interpretation of the negative-sample reinforcement term in wd1 as a form of data unlearning. Under the DCE likelihood approximation, the negative-sample reinforcement term in wd1 becomes
$\mathbb{E}_{o\sim p^{\prime}_{0}(\cdot)}\bigg[w^{-}(q,o)\cdot\log\pi_{\theta}(o|q)\bigg]=\mathbb{E}_{o\sim p^{\prime}_{0}(\cdot)}\bigg[\exp(-A(x_{0}))\cdot\log\pi_{\theta}(o|q)\bigg]$ (67)
$=$ $\mathbb{E}_{x_{0}\sim p^{\prime}_{0}(\cdot)}\bigg[\exp\big(-A(x_{0})\big)\cdot\underbrace{\mathbb{E}_{t\sim[0,T],p^{\prime}_{t|0}(x_{t}|x_{0})}\big[\sum_{x^{i}_{t}=[\texttt{mask}]}-\frac{1}{t}\log p_{\theta}(x_{0}^{i}|x^{\text{UM}}_{t})\big]}_{\mathcal{L}_{\text{DCE}}}\bigg]$ (68)
$=$ $\mathbb{E}_{x^{-}_{0}\sim p_{\text{data}}}\bigg[\underbrace{\mathbb{E}_{t\sim[0,T],p^{\prime}_{t|0}(x_{t}|x^{-}_{0})}\big[\sum_{x^{i}_{t}=[\texttt{mask}]}-\frac{1}{t}\log p_{\theta}(x_{0}^{-,i}|x^{\text{UM}}_{t})\big]}_{\mathcal{L}_{\text{DCE}}\ \Leftrightarrow\ \text{ELBO}}\bigg],$ (69)
where $p_{\text{data}}(x^{-}_{0})=p^{\prime}_{0}(x^{-}_{0})\frac{\exp(-A(x^{-}_{0}))}{\sum_{x^{-}_{0}}p^{\prime}_{0}(x^{-}_{0})\exp(-A(x^{-}_{0}))}$. Equation˜69 holds by simply applying importance sampling.
Since DCE is equivalent to the evidence lower bound (ELBO) of masked discrete diffusion models, we draw an analogy between the final objective in Equation˜69 and data unlearning in diffusion models (Alberti et al., 2025). Equation˜69 can be viewed as a direct masked discrete–diffusion extension of NegGrad (Golatkar et al., 2020), which aims to minimize the evidence lower bound of the log-likelihood on samples with lower advantage.